OpenStudy (anonymous):

A company sells boxes of bracelets (b) for $30 and boxes of necklaces (n) for $50. They make batches of bracelets that fill 5 boxes and batches of necklaces that fill 7 boxes. The company only has 48 boxes. Which system of equations helps the company plan to make $320? b = 48 - n 5b + 7n = 48 b = 48 - n 30b + 50n = 320 5b = 48 - 7n 30b + 50n = 320 5b = 48 - 7n 30b - 50n = 320

OpenStudy (anonymous):

b = 48 - n 5b + 7n = 48 b = 48 - n 30b + 50n = 320 5b = 48 - 7n 30b + 50n = 320** this one? 5b = 48 - 7n 30b - 50n = 320

OpenStudy (anonymous):

They're in sets of two

OpenStudy (anonymous):

5b + 7n = 48 -->5b = 48 - 7n 30b + 50n = 320 I agree with you :)

OpenStudy (anonymous):

I took an educated guess! :D yay, thanks! I have a few more word problems.. mind helping?

OpenStudy (anonymous):

Can't make no promises, but I will see what I can do :)

OpenStudy (anonymous):

Two competing internet companies have recently restructured their payment plans. Quick Net requires a $10 activation fee and charges $55 per month for their basic plan. Speedy Net requires no activation fee and charges $57 per month for their basic plan. When will the two plans be equal in cost? At 4 months At 2 months At 5 months (this) They will not ever be equal

OpenStudy (anonymous):

10 + 55m = 57m 10 = 57m - 55m 10 = 2m 5 = m answer is 5 months

OpenStudy (anonymous):

^_^.. Captain Liz is going to load her cruise ship with bags of sugar. She has room for 19 bags of sugar. She has allotted for 121 pounds of sugar on her ship. She can choose from 7-pound bags (the variable s) or 10-pound bags (the variable p). Which equations would be a constraint to Captain Liz maximizing her sugar load? 7s = 121 + 10p 7s = 121 - 10p s - p = 19 (this?) s = 19 + p

OpenStudy (anonymous):

or the last one

OpenStudy (anonymous):

s + p = 19 -- > s = 19 - p 7s + 10p = 121 --> 7s = 121 - 10p Your question is asking which EQUATIONS, does that mean there is more then 1 answer ?

OpenStudy (anonymous):

Sorry, internet problems.... **No, there is only one answer When planning his crops, Farmer Joe knows that his 5 acres can support corn and strawberries. He wants to make $5,500 from his crops. He can make $1,000 per acre of corn (the variable c) and $1,500 per acre of strawberries (the variable s). Which equation below would be a constraint in his system of equations? c + s = 5 (this) c - s = 5 1,500c + 1,000s = 5,500 1,500s - 1,000c = 5,500

OpenStudy (anonymous):

@kelliegirl33 my OS is messed up right now

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

I think the question before is : 7s = 121 - 10p and I agree with you on the other problem... c + s = 5

OpenStudy (anonymous):

for the previous question, I was going to say s + p = 19, but that is not one of the answers

OpenStudy (anonymous):

I thought that as well, okay. OS won't let me open a new question so.. When analyzing three linear functions, Bill noticed that they all intersected at the point (10, 10). What conclusion can Bill come to about the linear function with the largest slope? It has the highest y-intercept. (this maybe?) It has the middle y-intercept. It has the lowest y-intercept. Bill cannot conclude anything.

OpenStudy (anonymous):

sorry...I do not know this one....I am not good with functions :(